### v1

# High-speed ditching of double curvature specimens with cavitation and ventilation

Abstract
The water entry at high horizontal speed of double-curvature specimens, reproducing the rear part of the fuselage that first gets in contact with the water during aircraft ditching, is investigated experimentally. Three shapes are analysed, representing different aircraft types. Pressure and load measurements are taken, supported by underwater high-speed visualization. The effects of different horizontal speeds, pitch angles, and curvatures are analysed, while keeping a constant vertical-to-horizontal velocity ratio. It is observed that the longitudinal curvature plays a significant role in the hydrodynamics, potentially leading to cavitation and ventilation in the rear part of the specimen at high speeds. The transverse curvature affects the pressures and loads both at the front and at the rear, since a lower transverse curvature increases the possibility of the fluid to escape from the sides. An increase in pitch angle results in an increased loading, in terms of both amplitude and distribution. Finally, an increased horizontal speed leads to higher loads at the front but also affects the cavitation and ventilation modalities at the rear. The load scaling for flat plates also applies to these specimens at the front, but the occurrence of cavitation and ventilation invalidates it at the rear. The time evolution of the cavitation region and of the wetted area are examined through advanced image processing techniques and pressure data analysis. The comparison with the geometric intersection area of the specimen with the still water surface provides further insight into the hydrodynamic phenomena during the water entry.
[…]
1 Introduction
Aircraft ditching is the manoeuvre performed when the pilots and crew are forced to recur to an emergency water landing due to exceptional circumstances such as engine failure, fuel exhaustion, severe weather or other critical and unexpected situations. The goal of ditching is to land the aircraft safely on the water surface, allowing the occupants to be evacuated and rescued with the lowest risk of injury. Although being a rare event, airworthiness regulations require a precise certification for ditching, including demonstration tests proving that the aircraft structural damage during this eventuality is limited and that a sufficient floating time is guaranteed for a safe evacuation of the occupants (Climent et al., 2006).
The numerical approach for certification is becoming feasible but is is still in a developing phase (Siemann & Langrand, 2017; Siemann et al., 2018; Streckwall et al., 2007; Guo et al., 2013; Hughes et al., 2013; Bisagni & Pigazzini, 2018; Anghileri et al., 2011, 2014; Xiao et al., 2017), mainly because of the high resolution needed to correctly capture the hydrodynamic and the fluid-structure interaction phenomena, which implies quite large computational efforts. In order to ascertain the accuracy of the computational models and the reliability of the results, a careful validation and verification is deemed strongly necessary.
The experimental investigation allows to correctly reproduce the physical phenomena occurring during a ditching event. However, ditching tests of a full aircraft are impractical and also too expensive. Attempts to perform experiments mimicking the ditching scenario have been made during the last decades, starting from the pioneering work of Smiley (1951) who performed tests on flat plates entering water with a relevant horizontal speed. Smiley was able to retrieve accurate measurements of pressures, loads and of the wetted area at different test conditions. Even if remarkable for that time, this work has important limitations especially in terms of the data sampling frequency which, being of order of 1 kHz, does not enable to capture the fast transients occurring in a water entry problem (Van Nuffel et al., 2013). Another notable experimental ditching campaign was performed by McBride & Fisher (1953), who conducted free fall tests of several scaled fuselage models, representative of different types of aircraft. The study mainly focused on the role of the longitudinal and transverse body curvatures on the dynamics of the aircraft during ditching. The main limitation of the study lies in the impossibility in the control the impact conditions. In addition, information on the hydrodynamics developing beneath the models were provided, where it is expected that suction forces and possibly even ventilation phenomena may develop, as shown in Tassin et al. (2013).
Scaled model test of a full aircraft model were performed in Climent et al. (2006) and in Zhang et al. (2012). However, the commonly used scaling based on the Froude approximation does not apply in this case, in which highly non-linear hydrodynamic phenomena take place. Furthermore, the aircraft models used in the tests were too rigid, and therefore the fluid-structure interaction aspect could not be correctly reproduced. Focusing on hydrodynamic aspects only, since viscous effects are known to have a negligible role during water entry (Abrate, 2011; Facci et al., 2015; Moghisi & Squire, 1981), the similarity in terms the ventilation and cavitation numbers (Brennen, 2014) is deemed very critical. Such similarity conditions could be achieved only by performing tests in a de-pressurized environment.
In order to overcome such limitations, a High Speed Ditching Facility (HSDF) was built at the CNR - Institute of Marine Engineering. The HSDF allows tests at horizontal speed up to 48 m/s, vertical to horizontal velocity ratios ranging from 0.03 to 0.05 and on test models about 0.7 m wide and 1.3 m long. The pitch angle of the impacting body can be varied from 4 to 10. Depending on the needs, pressures, strains and forces acting on the body can be measured through a ruggedized acquisition system, hosted inside an acquisition box firmly connected to body. High-speed underwater videos are also recorded to help the interpretation of the measurements and to provide additional insights about the physical phenomena taking place during the impact (Iafrati et al., 2015).
Early tests were performed on a flat aluminium plates, 15 mm thick, in a configuration similar to that used in Smiley (1951). Different test conditions were examined by varying the horizontal speed, the vertical-to-horizontal velocity ratio and the pitch angle. The strong correlation between the spray root propagation velocity and the intensity of the pressure peak was confirmed and a scaling of the total loading with respect to the impacting velocity relative to the body was observed (Iafrati, 2015, 2016). Significant fluid-structure interaction phenomena were observed by carrying out tests of thinner aluminium plates, 3 mm and 0.8 mm thick. Differently from the expectations, it was found that higher hydrodynamic loads, up to 30%, occur in presence of large deformations of the plates and, therefore, the estimate of the loads based on tests on rigid specimens is not always conservative (Spinosa & Iafrati, 2021).
Although the tests on flat plates provided a wide dataset for the validation of the capabilities of the computational models to describe the hydrodynamics and the strong fluid-structure interaction phenomena occurring during ditching, the lower portion of an aircraft fuselage is characterised by a curved shape. In order to assess the effects of the transverse curvature of the fuselage, tests on both concave and convex plates, still 15 mm thick, were performed (Iafrati, 2018). However, the results of some numerical simulations and experimental tests on full fuselages showed that the longitudinal curvature at the rear may generate suction forces, associated with negative pressures, which affect the aircraft dynamics during the ditching (Climent et al., 2006; Zhang et al., 2012). In order to better investigate the phenomenon, an experimental test campaign on double curvature rigid specimens was conducted. From a preliminary analysis of the data it was found that cavitation and ventilation may occur at high speed and they may have important effects on the magnitude and distribution of the loading (Iafrati & Grizzi, 2019; Iafrati et al., 2020).
In the present work, a deeper analysis of the tests performed on double curvature specimens is provided. The objective of the analysis is to investigate the role played by the various parameters of the tests, specifically horizontal velocity, pitch angle and body curvatures. The analysis is presented in terms of pressures and total loads acting on the impacting specimens. The availability of the data of the loads at both rear and front makes it possible to identify the significant changes in the load distribution, when cavitation and ventilation phenomena occur. The evolution in time of the wetted area and of the cavitation zone is also investigated based on underwater visualisations, and quantitative estimates are also provided.
2 Experimental Set-up
2.1 The High-Speed Ditching Facility
The experiments are performed in the High Speed Ditching Facility (HSDF) of the CNR - Institute of Marine Engineering, described in detail in Iafrati et al. (2015), in which the test repeatability is also assessed. For the purpose of the present paper, it is worth recalling the most relevant features.
The facility is composed of a guide along which a properly designed trolley can freely move. The guide is about 64 m long and it is suspended by five bridges over the CNR-INM towing tank. The guide can be set at different inclinations with respect to the still water level of the basin, thus yielding different vertical-to-horizontal velocity ratios, in the range of 0.03 to 0.05.
The trolley carries a waterproof box that hosts a ruggedized case, with the acquisition system inside. The specimens to be tested are directly connected to the the bottom of the trolley. The trolley is accelerated by a system of eight bungee cords, set in a V-shaped configuration, but it is left to move freely along the guide shortly before the impact of the specimen with the water. Hence, the hydrodynamic loads acting on the model and the constraint forces that the trolley wheels exert on the guide are the only forces acting on the system, since gravity can be considered negligible. The total mass of the trolley and acquisition box is about 900 kg, which is large enough to ensure that the horizontal velocity remains nearly constant during the impact, being the velocity reduction during the impact phase lower than 5% of the initial velocity.
A video showing the execution of one of the experiments can be found at the link https://www.youtube.com/watch?v=1j5-Z96NNyQ.
2.2 Specimen shapes
The tested models are double curvature specimens that reproduce the part of the fuselage body that first gets in contact with water during ditching. Just to give an idea, the typical shape of the specimen is the yellow portion of the entire fuselage shown in Figure 1.
(a) Side View
(b) 3D View
Figure 1: Drawing of the whole fuselage S2 (blue) and of the extracted specimen part (yellow) in (a) side view and (b) 3D view. The local coordinate system $(\bar{x},\bar{y},\bar{z})$ is also shown.
Three fuselage shapes are considered, denoted here as S1B, S2 and S3, which are described in a fully analytical form in Iafrati et al. (2020). For the readers’ convenience the details of the analytical formulation are also provided in the following.
The shapes are defined in terms of a few non-dimensional parameters, namely, Length-to-Breadth ratio ($L_{B}$), Front length-to-Breadth ratio ($F_{B}$) and Rear length-to-Breadth ratio ($R_{B}$), where $B$, the maximum breath of the fuselage, is assumed as the reference length value. Hence, non-dimensional coordinates are introduced as: $\xi=\bar{x}/B$, $\eta=\bar{y}/B$ and $\zeta=\bar{z}/B$ (Figure 2). All shapes are defined by a central region with a constant circular cross-section: a fore part where the circular cross-sections are uniformly and progressively reduced about the centre of the circle until the nose, and a rear part where the shape is shrunk, with the centre of the circle shifted in the longitudinal plane in a way that the upper points are always located at $\zeta=0.5$.
(a) Midplane section of S1B
(b) Midplane section of S2 and S3
(c) Circular-elliptical cross-section of S3
Figure 2: Non-dimensional mid-plane sections of the fuselages S1B (a) and S2/S3 (b), where $\xi_{i}=\bar{x}_{i}/B$ and $\xi_{e}=\bar{x}_{e}/B$, see Table 1. In (c) the construction of the semi-elliptical cross-section of the shape S3 is sketched.
The shapes S1B and S2 have both a circular cross-section whose radius $\hat{r}(\xi)$ varies with the longitudinal coordinate $\xi$ as follows:
$\hat{r}(\xi)=\left\{\begin{array}[]{c}0.5\sqrt{1-{\left(\frac{\xi-F_{B}}{F_{B}}\right)}^{2}}\qquad 0<\xi<F_{B}\\ \\ 0.5\qquad F_{B}<\xi<\left(L_{B}-R_{B}\right)\\ \\ 0.5+\hat{o}_{C}(\xi)\qquad\left(L_{B}-R_{B}\right)<\xi<L_{B}\end{array}\right.$ (1)
where the offset function $\hat{o}_{c}(x)$ is defined as
$\hat{o}_{C}(\xi)=-0.5\sin{\left(\frac{\xi-(L_{B}-R_{B})}{K\,R_{B}}\right)}\cdot\left[\frac{\xi-(L_{B}-R_{B})}{\sin{\left(1/i\right)}\,R_{B}}\right]^{1/i}$ (2)
$i$ being a non-dimensional parameter. The centre of the cross-section is located at $\textbf{{O}}\equiv(\xi,0,0)$ for $\xi\leq(L_{B}-R_{B})$ and at $\textbf{{O}}\equiv\left(\xi,0,\hat{o}_{C}(\xi)\right)$ for $\xi>(L_{B}-R_{B})$, thus ensuring that the sections are all aligned at $\zeta=0.5$ in the rear fuselage portion.
The shape S3 has a circular cross-section blended with an elliptical cross-section at the bottom. Starting from the circular profile, the elliptical portion is defined by two parameters, i.e the angle $\theta$ and the ratio $c/h$ of the shorter ellipse semi-axes to the offset $h$, see Figure 2(c). The contact and the tangency between the circumference and the ellipse are enforced at the contact points $M$ and $N$, by choosing the two semi-axes $a$ and $c$ as:
$a=\frac{c}{h}\,\frac{\sin{\theta}}{\sqrt{{\left(\frac{c}{h}\right)}^{2}-1}}\qquad c=\frac{a\,{\tan{\theta}}}{\sqrt{{\left(\frac{c}{h}\right)}^{2}-1}},$ (3)
and therefore the bottom part of the section is described by the equation
$\left(\frac{\eta}{a}\right)^{2}+\left(\frac{\zeta+1-(e+c)}{c}\right)^{2}=1,$ (4)
where
$e=1-\cos{\theta}-c\,\left(1-h/c\right)\;.$ (5)
When the circular-elliptical section is used, the equation of the offset is modified as
$\hat{o}_{\textrm{CE}}(\xi)=2\hat{o}_{C}(\xi)/e.$ (6)
which ensure that the bottom profiles of shape S2 and S3 are perfectly overlapped.
The parameters that define the three fuselage shapes are listed in Table 1. Therein, the dimensional parameters $\bar{x}_{I}$ and $\bar{x}_{E}$ defining the limits of the specimen in the longitudinal direction are also provided. Note that in all cases the specimens are 1.24 m long and 0.66 m wide.
Parameter S1B S2 S3
$L_{B}$ [-] 8 7.5 7.5
$F_{B}$ [-] 1.5 1.5 1.5
$R_{B}$ [-] 2 2.5 2.5
$K$ [-] 2 1.55 2.55
$i$ [-] 1 2.6 2.6
$C/H$ [-] - - 5
$\theta$ [] - - 50
$B$ (m) 1 1.5 1.5
$\bar{x}_{I}$ (m) 5.7 6.71 6.71
$\bar{x}_{E}$ (m) 6.94 7.95 7.95
Table 1: Parameters defining the whole fuselage shapes and the extracted specimens
It is worth noting that the specimen S1B is extracted from a fuselage with a breadth (diameter) of 1 m, whereas the fuselage S2 and S3 from a fuselage with a breadth (diameter) of 1.5 m. The different choice of the was motivated by the need of maximizing, in the case of the specimen S1B, the vertical distance between the lowest point, i.e. the first contact point, and the leading edge of the specimen, which defines the time at which the front of the specimen gets below the still water level and thus the end of the impact phase.
The difference among the three shapes, set at a pitch angle of 6 , is highlighted in 3(c), in which a three-dimensional view is shown, whereas the longitudinal and transverse sections passing through the point of first contact with water, are shown on 3(a) and (b) respectively.
Figure 3: Comparison among the geometry of the three specimens aligned with respect to the lowest Z point, i.e the point of first contact with water: (a) section on the longitudinal plane (b) section on the transverse plane and (c) 3D view
For the experimental data analysis two reference frames are defined: a ground fixed reference frame and a trolley-fixed (or specimen-fixed) reference frame, both shown in Figure 4.
Figure 4: Reference frames, coordinate systems and points of interest to study the water entry of the shapes S1B (a) and S2/S3 (b) at a pitch angle of $\alpha=6^{\circ}$.
The coordinates associated with the ground-fixed reference frame are indicated with upper-case letters $X$, $Y$ and $Z$. In particular, $X$ is parallel to the undisturbed free surface, $Z$ is vertical and directed upwards and $Y$ is in the direction on a right-handed coordinate system. The specimen horizontal and vertical speed are defined in this reference frame as $U=dX/dt$ and $V=-dZ/dt$, respectively. The coordinates associated with the trolley-fixed reference frame are indicated with lower-case letters $x$, $y$ and $z$, with origin O. As shown in Figure 4 the origin is located at the leading edge of the plate. The $x$ coordinate is parallel to the longitudinal axis of the trolley, the $z$-coordinate is normal to it and directed upwards, whereas the $y$ axis forms and right-handed coordinate system. The pitch angle $\alpha$ is defined as the angle between the trolley axis $x$ and the $X$-direction, whereas the angle $\beta=\arctan{V/U}$ is introduced to denote the angle between the the impact trajectory of the specimen and the undisturbed water level.
For each shape, at given $\alpha$ and $\beta$, it is worth defining the point C as that where the longitudinal curvature starts changing at the rear, which is the point with dimensional $x$-coordinate $(L_{B}-R_{B})$ in Figure 2(a,b), and the point H, which defines the lowest point of the specimen, and thus the first getting in contact with water. As it is possible to see, the two points are very close to each other for the fuselage shapes S2 and S3 and relatively far in case of shape S1B.
2.3 On-board Measurements
Pressures are acquired through thirty pressure probes type Kulite XTL 123B, full scale range of 300 psi, properly calibrated before starting the test campaign. The probes are installed flush to the external surface of the specimen with $\pm$ 0.1 mm accuracy. The position of the probes is given in Figure 5. In the analysis of the data, beside the specific probe number, reference is also made to probe lines ($l1,l2,l3,l4$) and rows ($r1,r2,...,r8$) as indicated in the figures.
(a) Shape S1B
(b) Shapes S2 and S3
Figure 5: Position of the pressure probes and definition of the row and lines used for the discussion. Dimensions are in mm.
The pressure time histories presented in the following are raw data, unless otherwise specified. It is worth noting that sometimes pressure spikes of non-physical origin may appear, which, among other reasons, are due either to slightly protruding probes generating at such speeds cavitating vortices (Iafrati et al., 2015) or to the temperature shock experienced by the sensors when they get in touch with water (Van Nuffel et al., 2013), which is at a lower temperature than air.
The acquisition box hosting the specimen is connected to the trolley by a total of six single-axis load cells, installed between two floating pins to minimize the loading in other directions, as shown in Figure 6.
Figure 6: Position of the load cells, referring to the specimen S2. The total rear and front force direction is indicated, as well as the trolley-fixed reference frame. The dimensions are in mm. The value of $H$ is 520 mm, 440 mm and 436.2 mm for the shapes S1B, S2 and S3, respectively.
Four Kistler 9343 cells, with full scale range of 70 kN, are used in the $z$-direction, whereas two Kistler 9363 cells, with full scale range of 120 kN, are used in the $x$-direction. The load cell signals are conditioned using a charge amplifier type Kistler 5073-A421. The channels are connected in a way that the loads acting in the $z$-direction can be distinguished in the rear $F_{zR}$ and in the front $F_{zF}$ contributions, whereas $F_{x}$ denotes the load acting in the $x$-direction. The distance $H$ for the three specimens S1B, S2 and S3 are 520 mm, 440 mm and 436.2 mm respectively. Although the single load cell is very accurate when loaded directly along its axis, the connection to the frame of the acquisition box introduces some uncertainties, since the loading is not exactly aligned with the axes of the cells. Furthermore, due to the inherent free play of the installations, some errors may be expected for low values of the loading. For this reason a careful calibration was carried out before starting the test campaign, and it was found that the recorded data are always below the actual loading. The underestimate ranges from about 8% for a load of order of 8 kN and reduces continuously when increasing the load, being about 4% for a load of order of 14 kN. The load was also applied at different locations along the $x$-axis and the centre of loads could be retrieved with a satisfactory accuracy by the combined data of the rear and the front load cells. Unless specified differently, also the force time histories provided in the following are raw data.
All the data are acquired by a ruggedized acquisition system hosted in the acquisition box. The acquisition system is composed of four Dewesoft© Sirius modules and one Dewe43 module. All channels are sampled at 200 kHz with a double level A/D 24 bit conversion.
2.4 High-Speed Cameras
Two high speed cameras are used to record the tests. One camera, type Photron FastCam Mini AX, operating at 5000 fps, is installed at the side of the facility at the location of the water entry, and it is mainly used to retrieve the horizontal velocity at the exact moment the specimen touches the water. A second camera, type Photron FastCam SA-X2, operating at 3000 fps, is installed on a properly designed waterproof case and positioned underwater. This camera records the water entry from below and provides important information for a more complete interpretation of the hydrodynamic phenomena and of the pressures/loads measurements. Both cameras are synchronized with the on-board acquisition systems.
2.5 Test Conditions
The test conditions that are discussed in the present paper are listed in Table 2. All tests presented here are performed guide inclination angle $\beta=\arctan(V/U)$. Being $V/U=1.5/40=0.0375$, then $\beta=2.1476^{\circ}$. It is worth noticing that, due to the use of bungee ropes to accelerate the trolley, a very precise control of the speed cannot be achieved, which is the reason why the actual values of the horizontal speed are not exactly the nominal expected values. It should be noted that in the discussion from now on, however, reference will be made to the nominal velocity values and not to the actual ones reported in the table.
The three groups refer to the test cases examined to investigate the effect of the horizontal speed, of the shape and of the pitch angle. The test condition highlighted in bold is the one used as a baseline test condition when examining the effects of the above-mentioned parameters.
Shape Test ID U V Pitch V/U $t_{0}$ $t_{E}$ $\left(t_{E}-t_{0}\right)$
Effect of speed
S2 1H2022_23_07_2018_1 29.87 1.12 6 0.0375 0.213265 0.28034 0.067075
S2 1H2122_19_07_2018_1 34.5 1.29 6 0.0375 0.222075 0.273589 0.051514
S2 1H2222_17_07_2018_1 40.25 1.51 6 0.0375 0.24621 0.295279 0.049069
S2 1H2322_18_07_2018_1 46.19 1.73 6 0.0375 0.241675 0.285315 0.04364
Effect of shape
S1B 1H4222_06_09_2018_1 40.45 1.52 6 0.0375 0.25908 0.286411 0.025965
S2 1H2222_17_07_2018_1 40.25 1.51 6 0.0375 0.24621 0.295279 0.049069
S3 1H3222_08_06_2018_1 40.25 1.51 6 0.0375 0.240705 0.287465 0.04676
Effect of pitch angle
S2 1H2221_17_07_2018_1 40.57 1.52 4 0.0375 0.252580 0.283938 0.03136
S2 1H2222_17_07_2018_1 40.25 1.51 6 0.0375 0.24621 0.295279 0.049069
S2 1H2223_16_07_2018_1 40.08 1.50 8 0.0375 0.216710 0.283750 0.06704
Table 2: Test Conditions.
3 Results and discussion
3.1 Definition of the impact phase and of $\tau$
For the purpose of the present study, the data analysis is limited to the impact phase, which lasts from the time of first contact of the specimen with the water surface, $t_{0}$, up to the time $t_{E}$ at which the root of the forward propagating spray reaches the leading edge of the specimen. As done in Iafrati & Grizzi (2019), $t_{0}$ can be approximated as the time at which the first pressure probe exhibits a pressure growth, whereas $t_{E}$ is estimated by linear extrapolation starting from the times at which the pressure peak passes over the location of the probes P28 and P30 (Iafrati, 2016). As shown in Figure 4, for the shapes S2 and S3 at a pitch angle of 6 the estimation of $t_{0}$ based on the time history of the pressure measured by the probe P17 is expected to be quite accurate, being the probe P17 located very close to the point of first contact H. Instead, for the shape S1B the first contact point is located between P21 and P24, therefore the estimate of $t_{0}$ is more uncertain. The estimate based on pressure measurements is corrected by computing the time duration of the impact phase from the underwater videos and subtracting it from $t_{E}$.
Once $t_{0}$ and $t_{E}$ are defined, a non-dimensional time variable, denoted as $\tau$ can be defined as
$\tau\doteq\frac{t-t_{0}}{t_{E}-t_{0}}\;.$ (7)
The use of the non-dimensional time $\tau$ allows a fairer comparison of the results obtained at different test conditions, and, in particular, at different pitch angles.
3.2 Hydrodynamics phenomena during the water entry
Qualitative considerations on the hydrodynamics of the water entry of the double curvature specimens can be made through a comparative analysis of the pressure time histories and the underwater high-speed videos. Such an analysis is conducted on the test of the shape S2, advancing with a horizontal velocity $U$=40 m/s, velocity ratio $V/U$=0.0375 and pitch angle $\alpha$= 6. Video frames taken at different $\tau$, displaying the flow evolution on the bottom of the specimen, are shown in Figure 7.
(a) $\tau$ =0
(b) $\tau$ =0.1
(c) $\tau$ =0.2
(d) $\tau$ =0.3
(e) $\tau$ =0.4
(f) $\tau$ =0.5
(g) $\tau$ =0.6
(h) $\tau$ =0.7
(i) $\tau$ =0.8
(j) $\tau$ =0.9
Figure 7: Frames of the underwater video at different $\tau$ during the water entry of the shape S2 at a nominal horizontal speed of $U$=40 m/s and at a pitch angle of $\alpha$=6. The specimen moves from left to right.
The time histories of pressure measured by probes located in the same row (see Figure 5 for the row definitions) are shown in Figure 8.
(a) Probe positions: the point C separates the part of the specimen with zero curvature on the right, from that with finite curvature on the left; H is the point of first contact at $\alpha$=6.
(b) Row $r8$
(c) Row $r7$
(d) Row $r6$
(e) Row $r5$
(f) Row $r4$
(g) Row $r3$
(h) Row $r2$
(i) Row $r1$
Figure 8: Pressure probe time histories row by row for the shape S2 moving at $U$=40 m/s, at a pitch angle of $\alpha$=6.
From the underwater frames shown in Figure 7 it is possible to recognize the wetted area as the darker region. During the early stage of impact, i.e. in Figure 7$(b,c,d)$, the wetted area grows in all directions. The brighter zone around the wetted area represents the spray region.
In the front part of the specimen the spray propagates in a way similar to that observed during the water entry of a curved plate, as shown in Iafrati (2018). The spray root is characterized by a localized pressure peak that causes a sharp rise of the pressure recorded by the probes, as shown in Figure 8 for rows from $r5$ to $r8$. After the passage of the peak, the pressure gradually diminishes, similarly to what found for the flat plates (Iafrati, 2016; Iafrati et al., 2015) or convex plates (Iafrati, 2018). As observed in Iafrati et al. (2015), the pressure drop recorded by some of the probes (see for example P22, P25 or P26) before the sharp pressure rise is associated with the interaction of the flow in the spray tip with the probe itself, which can also induce cavitating vortices. For the probes located at the row $r5$, which for the shape S2 at 6 pitch is just ahead of the point of first contact, all probes approach a constant negative pressure value, slightly higher than the vapour pressure. From both the underwater images it is possible to see that the spray root is curved backwards, due of the combined action of three-dimensional effects and of the transverse curvature of the specimen, as observed in flat plates (Iafrati, 2016), and in convex plates (Iafrati, 2018), respectively. This is also confirmed by the time histories of the pressures, from which it is observed that the spray root region, i.e. the pressure peak, reaches first the probes located in the mid-line, such as P21, P24, P28 and P30, and next the probes located at the sides, see for instance the recordings of the probes on $r5$ in 8(e) . In Figure 8(d), the time histories of probes P23 and P25, are very close to each other, with the pressure peak exhibiting about the same delay compared to the probe P24, confirming that a satisfactory level of symmetry is achieved.
In the rear part of the specimen, i.e. the region behind C, the hydrodynamics is quite different from that observed in the front part. In the early stage the wetted area propagates backwards, although at a lower speed than that observed in the front part. This is coherent with the lower the values of the pressure peaks measured by the probes located at the rear, such as P17. Soon after the passage of the spray root, the pressures decrease quite rapidly and become negative. Such a behaviour is associated to the shape of the body and with the impact conditions. A qualitative explanation can be provided from two different points of view, i.e. focusing on the flow in the midplane (longitudinal plane) or on the flow in the transverse plane.
By focusing on the longitudinal plane (midplane) in the trolley-fixed reference frame, the results of the two-dimensional case of an asymmetrical wedge entering water with a horizontal velocity (Judge et al., 2004; Krastev et al., 2018) can provide some elements for the interpretation of the flow phenomena in the present case. In fact, as a consequence of the kinematic asymmetry of the flow due to the horizontal speed, the pressure on the front part of the specimen is higher than that at the rear. From a theoretical standpoint, it is expected that a velocity singularity will occur at the apex (Judge et al., 2004). In real flows, a strong local velocity acceleration is observed there, accompanied by a pressure drop, which is similar to what occurs on the suction side of hydrofoils (Faltinsen, 2005; Newman, 1971). The flow field presents also similarities with that of the vertical water entry of inclined asymmetric wedges (Semenov & Iafrati, 2006; Riccardi & Iafrati, 2004). In Xu et al. (2008), in which a wedge entering water with both a vertical and a horizontal velocity is studied using a self-similar approach, even if the velocity ratio $U/V$ is much lower than the present case, a local pressure drop to negative values is observed at the wedge apex.
An alternative interpretation of the pressure behaviour can be obtained by focusing on the flow in the transverse plane, in an earth-fixed reference frame, by exploiting the so called 2D+$t$ approach (Fontaine & Tulin, 1998; Iafrati & Broglia, 2010). In the front part the plate always penetrates into the water, whereas in the rear part, depending on the local inclination of the tangent to the body surface in the longitudinal plane and on the $U/V$ velocity ratio, the fluid flow problem changes from a water entry to a water exit one (Tassin et al., 2013) and when this happens, negative pressures develop. The development of low and even negative pressures occurs very shortly after the first contact with water and, based on the above considerations, the pressure drop is expected to be higher if the gradient in the slope of the body surface is more significant, which means in presence of larger longitudinal curvature.
The simulations of the horizontal planing motion of the same scaled fuselage models at low speeds and at the same pitch angle, discussed in Spinosa et al. (2022), may also offer insight on the hydrodynamics of the water entry at at high horizontal speed. This work also demonstrates the attainment of negative pressure behind point H. Furthermore, it is shown that the pressure does not become negative abruptly at point H. Instead, the low pressure region extends slightly ahead of this point.
As observed in Iafrati & Grizzi (2019), the pressure drop can be so high that the vapour pressure value may be reached. Given the water temperature of the tests between 15 and 18  C, the vapour pressure relative to the atmospheric value, denoted as $p_{\textrm{vap}}$, is in the range -99.6 and-99.2 kPa. This explain the formation of a cavitation region (Acosta, 1973; Brennen, 2014), as shown in the underwater images of Figure 7 at the non-dimensional time $\tau\geq 0.2$. In the trolley-fixed reference frame, the cavitation region has a sort of reverse ”D” shape, which follows the pressure distribution on the bottom of the specimen. As the specimen enters in water, the forward front of the cavitation region maintains rather straight and located about the point of curvature change C, whereas the rear edge of the cavitation region spreads backwards towards the trailing edge of the specimen. Once the cavitation region reaches the trailing edge of the specimen, the water vapour inside it, at vapour pressure, comes into contact with air at atmospheric pressure. As a consequence, air is suddenly entrained into the water vapour cavity, forming a ventilation zone that expands in the opposite direction, i.e towards the leading edge. This behaviour is typical of the so called cavitation-ventilation regime, described in more detail in Section 3.3 based on the analysis of the water entry of the specimen S3 (Iafrati & Grizzi, 2019; Iafrati et al., 2020) and also occurs for S2 at the same horizontal speed.
The expansion of the cavitation region can be also recognized by observing the time histories of the pressure probes, see Figure 8. The passage of the rear edge of the cavitation region corresponds to the abrupt drops to the vapour pressure value. Moving backwards, i.e. from row $r4$ to row $r1$, the instant at which the drops occur for the midline probes (P17, P13, P9 and P4) increases. This time delay is related to the propagation velocity of the rear edge of the cavitation region. Comparing the pressure time histories across the same rows, it can be noted that pressure drops of the midline probes are delayed with respect to those of the probes at the sides, due to the combined effect of of the body shape and of the curvature of the rear edge of the cavitation region. When ventilation starts and air is entrained within the cavity, pressures start increasing progressively, reaching the zero-pressure value. It is worth noting that, in contrast to the pressure drop associated with the propagation of the cavitation region, the rise in pressure values during this phase is gradual and occurs at extremely close time instants for all pressure probes positioned on the same row. The increase in pressure values occurs later in time, when progressing from row $r1$ to $r4$. This clearly indicates that the ventilation front is advancing at an extremely high speed, as previously observed for the shape S3 in Iafrati & Grizzi (2019).
A comprehensive global overview of the shape and evolution of the cavitation region can be derived by constructing pressure iso-lines using the available pressure data from row $r1$ to $r5$ at different time instants. In Figure 9 the iso-lines overlapped to the corresponding video frames at $\tau$=0.3, 0.6 and 0.8 are shown.
(a) $\tau$= 0.3
(b) $\tau$= 0.6
(c) $\tau$= 0.8
Figure 9: Pressure iso-lines in the rear part of the specimen overlapped to the underwater video frames for the test of shape S2 at U=40 m/s, at $\alpha=6^{\circ}$ and at $\tau$=0.3, 0.6 and 0.8. The symmetry about the mid-line is exploited.
Although the probe spacing is not fine enough to provide a very accurate interpolation, especially in this case where relevant pressure gradients are present, it is evident that after the first contact with water, the pressure iso-lines also assume the shape of a reverse D and expand in time in the interval between $\tau$= 0.3 and $\tau$= 0.6. The pressure values in time decreases overall and from a certain $\tau$ reach the value of $p_{\textrm{vap}}$ within the inner iso-line, which matches the shape of the cavitation region visible in the video frames. As the cavitation region reaches the trailing edge, the ventilation front starts moving back very fast towards the leading edge. The ventilation front is visible especially at $\tau$=0.8, where the iso-lines at p $\approx$ 0 have advanced towards the leading edge, having almost reached the point C. The ventilation front at this time instant appears to be quite straight in the transverse direction.
Observing the pressure time histories, it can be observed that ventilation occurs within the time interval between $\tau\approx 0.7$ and 0.8. By focusing the pressure iso-lines over this time interval, as shown in Figure 10, it is possible to gain more insight on the ventilation dynamics.
(a) Shape S2 - $\tau$=0.70
(b) Shape S2 - $\tau$=0.72
(c) Shape S2 - $\tau$=0.74
(d) Shape S2 - $\tau$=0.76
(e) Shape S2 - $\tau$=0.78
(f) Shape S2 - $\tau$=0.8
[…]
In fact, it is observed that the pressure begins to rise back to the atmospheric value first at the centre, since the cavitation region reaches the trailing edge first there. However, the ventilation front develops rapidly and at about $\tau$=0.78. becomes roughly parallel to the transverse direction, as also observed for the shape S3 in Iafrati & Grizzi (2019).
Given the considerations made in this section, based on the observation of the underwater video-frames and on the pressure time histories, a schematic showing the evolution of the cavitation region and the formation of the ventilation zone in the mid-plane is proposed in Figure 11.
Figure 11: Qualitative schematic describing the evolution of the cavitation region and of the ventilation zone in the mid-plane.
In Figure 12 the normal loads acting on the specimen recorded by the rear and front cells are plotted. The data of two test repeats, denoted as R1 and R2, performed at the same conditions are drawn. The red and blue vertical lines indicate the non-dimensional time at which the cavitation and the ventilation phases start, respectively. These time instants are determined by using the data of the pressure probes: the cavitation phase is assumed to start when the pressure at probe P17 drops to the vapour pressure value, while the ventilation phase when the pressure at probe P4 begins to rise to atmospheric pressure.
Figure 12: Time histories of the forces recorded by the front (red) and rear (blue) load cells and of the total normal load (black), for shape S2 at 6 pitch angle and $U$=40 m/s. The vertical red and blue lines indicate the start of cavitation and of ventilation respectively. The solid and dashed lines refer to the first and second test repeat respectively (denoted as R1 and R2).
The time histories of the force components exhibit some oscillations, which are more evident when the loads are low. These oscillations are partly due to the vibrations of the trolley, as it moves along the guide, and partly to the free play at the joints. In spite of these oscillations, the data of the two repeats display a quite similar trend. The loads measured by the front cells are always positive and grow progressively as time elapses, which is consistent with the positive pressures recorded by the probes located from $r5$ to $r8$ (see Figure 6). The load measured by the cells at the rear is rather low at the beginning, which is presumably associated with the development of cavitation, starting slightly before $\tau=0.1$ at this specific test condition. As the cavitation region expands, the loads measured at the rear undergo a reduction, which is rather evident between $\tau\approx 0.4$ and $\tau\approx 0.6$. As soon as the ventilation starts, the loads at the rear start increasing and turn positive. Such growth is consistent with the increase in the pressure at the rear of the body, when the pressure passes from the vapour pressure value to zero (atmospheric pressure). The load at the rear remains almost constant after $\tau\approx 0.8$, whereas the total load continues to grow up to $\tau=1$, when the spray root has reached the leading edge. At this time the front face of the trolley starts penetrating into the water, circumstance that reduces the horizontal speed and then the loads.
In the next sections, comparisons are established between tests performed at different conditions, in order to evaluate the effect of the horizontal speed, pitch angle and body curvature on pressures and loads.
3.3 Effect of the horizontal speed
As previously noted in Iafrati & Grizzi (2019); Iafrati et al. (2020), significant alterations in the hydrodynamics of the water entry are observed when the horizontal speed is varied for the shape S3. In particular, at least four regimes were identified: a no-cavitation regime (for $U$ lower than about 27 m/s), in which no cavitation is observed; an incipient cavitation regime (between $U\approx$ 27 m/s and 35 m/s) in which a cavitation region expanding in time only up to a certain $\tau$ can be seen; an intermediate regime, (between $U\approx$ 35 m/s and 37 m/s) (denoted in Iafrati & Grizzi (2019) as Full Cavitation Bubble and Cavitation Alternate with Ventilation), in which the cavitation region reaches the trailing edge, followed by the formation of a local ventilation region, which does not expand in relevant way; finally a cavitation-ventilation regime ($U$ higher than 37 m/s), in which the ventilation region exhibits a successive fast expansion towards the leading edge of the specimen.
In the present work, the effect of the horizontal speed is analysed for the shape S2, entering water with a pitch angle of 6 and at velocity ratio of $V/U=0.0375$. The analysis is performed in terms of pressures and loads and through the observation of the underwater images taken at two non-dimensional times, i.e. $\tau=0.2$ and $\tau=0.6$. A comparison is presented between the time histories of pressures recorded by the probes located on the midline and on line $l3$. (Figure 6). In order to make the comparison easier, the probes lying along rows $r5-r8$ and $r1-r4$ are plotted in separate graphs. The results for $U$=30 m/s and 35 m/s are shown in Figure 13, whereas those for $U$=40 m/s and 45 m/s in Figure 14.
(a) $U$=30 m/s - Frame $\tau$ =0.2
(b) $U$=35 m/s - Frame $\tau$ =0.2
(c) $U$=30 m/s - Frame $\tau$ =0.6
(d) $U$=35 m/s - Frame $\tau$ =0.6
(e) $p$ front probes - midline - $U$= 30 m/s
(f) $p$ front - probes midline - $U$= 35 m/s
(g) $p$ front probes - line l3 $U$= 30 m/s
(h) $p$ front probes - line l3 $U$= 35 m/s
(i) $p$ midline - rear probes - $U$= 30 m/s
(j) $p$ midline - rear probes - $U$= 35 m/s
(k) $p$ line l3 $U$= 30 m/s
(l) $p$ line l3 $U$= 35 m/s
Figure 13: Front and rear pressure probe time histories and under-water video frames comparison for the tests of shape S2, at a pitch angle 6 and $U$=30 m/s and 35 m/s.
(a) $U$=40 m/s - Frame $\tau$ =0.2
(b) $U$=45 m/s - Frame $\tau$ =0.2
(c) $U$=40 m/s - Frame $\tau$ =0.6
(d) $U$=45 m/s - Frame $\tau$ =0.6
(e) $p$ front probes - midline - $U$= 40 m/s
(f) $p$ front - probes midline - $U$= 45 m/s
(g) $p$ front probes - line l3 $U$=40 m/s
(h) $p$ front probes - line l3 $U$=45 m/s
(i) $p$ midline - rear probes - $U$=40 m/s
(j) $p$ midline - rear probes - $U$=45 m/s
(k) $p$ line l3 $U$=40 m/s
(l) $p$ line l3 $U$= 45 m/s
Figure 14: Front and rear pressure probe time histories and under-water video frames comparison for the tests of shape S2, at a pitch angle 6 and $U$=40 m/s and 45 m/s.
As shown in Figures 13 and 14 the probes located in the front, i.e. ahead of the point H, are characterized by the typical sharp rise, followed by a gentle decay. It is worth noting that, for all probes, the non-dimensional time $\tau$ at which the pressure peak is attained is almost independent of the horizontal speed, whereas the peak value increases with the speed. As expected, the peaks recorded by the probes along the line $l3$ occur at a later time with respect to the peaks recorded by the probes at the midline in the same row. This is a consequence of the combined effect of the transverse curvature of the specimen and of the curvature of the spray root, which is associated with the possibility for the water to escape from the sides (Iafrati, 2016).
The analysis of the pressure measurements of the probes located behind the point H is less straightforward and the support of the underwater images is needed for a correct interpretation. In fact, looking at the underwater frames in Figures 13 and 14 it can be seen that at $\tau$=0.2 the cavitation region is clearly visible only at $U$=40 m/s and 45 m/s. At $\tau$=0.6 the cavitation region is visible at all speeds, however at $U$=40 m/s and 45 m/s it has already reached the trailing edge of the specimen, whereas at $U$=30 m/s and at $U$=35 m/s it has expanded less significantly.
(a) $U$=30 m/s $\tau$ =1
(b) $U$=35 m/s $\tau$ =1
(c) $U$=40 m/s $\tau$ =1
(d) $U$=45 m/s $\tau$ =1
Figure 15: Frames of the underwater video taken at $\tau$=1 for the shape S2, at a pitch angle of $\alpha$=6 and at the different speeds.
By looking at the signals recorded by the probes located at the rear, all of them exhibit a descent to the vapour pressure value. The start time of the descent is consistent across all speeds, while the descent itself becomes sharper as the horizontal speed increases. At $U$=40 m/s and 45 m/s, all probes, except for P15 and P17 (located on $r4$), display a rise towards the atmospheric value due, to the occurrence of ventilation, which is consistent with what shown by the underwater images. The rise starts earlier at higher speed. The probes P15 and P17, being located in the region where the longitudinal curvature is rather high, remains in the cavitation region and are not reached by the ventilation front.
To gather more insight on the cavitation and ventilation modalities, especially in the final part of the impact phase, in Figure 15, the video frames at $\tau$=1 at all speeds are shown. It is possible to see that $U$=30 m/s at $\tau$=1 the cavitation region has expanded with respect to $\tau$=0.6, see Figure 13(c), but its rear edge has not moved significantly backwards. Therefore, this case can be ascribed to the incipient cavitation regime, also discussed in S3 (Iafrati & Grizzi, 2019). At $U$=35 m/s the cavitation region has reached the trailing edge and a ventilation region appears, which however remains confined around the trailing edge area, similarly to what happens in the intermediate regime, discussed above. Finally, at $U$=40 m/s and $U$=45 m/s $\tau$ =1 the ventilation front has clearly moved back towards the leading edge, therefore in these cases a cavitation-ventilation regime is observed. As a final comment, it can be observed that the threshold values that delineate the various regimes in S2 are analogous to those of S3.
As observed in (Iafrati, 2016), when varying the horizontal speed, the pressure should scale as $\rho\left(U^{2}+V^{2}\right)\,\sin{(\alpha+\beta)}$. As shown in 16, the scaling holds for the probes located at the front, whereas for the probes located at the rear the occurrence of cavitation introduces a lower bound for the pressures, which makes the scaling no longer valid. This is the reason why model scale tests could be reliable only if performed in de-pressurized conditions, which ensure the similarity of the cavitation number.
Figure 16: Time histories of the pressure scaled with the quantity $\rho\,(U^{2}+V^{2})$ for the tests of shape S2, at pitch angle 6 and different horizontal speeds.
The force acting in the $z$-direction and measured by the load cells located at the front and at the rear are shown on the upper panels of Figure 17.
Figure 17: Time histories of the forces for the tests of shape S2, at pitch angle 6 and different horizontal speeds.
The normal force recorded by the front load cells $F_{zF}$ is more sensitive to the pressures acting on the front portion of the specimen, and for this reason it displays an increasing trend, which is consistent with the forward propagation of the spray root. The force intensity grows with speed, in line with the corresponding increase of the pressures. On the other hand, the rear load cells are more sensitive to the pressures acting on the back portion of the specimen and then, aside from the very early stage of the impact phase, they undergo the effects of the cavitation and ventilation phenomena. The load at the rear $F_{zR}$ exhibit a similar behaviour at the different speeds, assuming more and more negative values as the cavitation area expands up to about $\tau$= 0.5. Successively, when the ventilation starts, earlier for the test at $U$=45 m/s and then at 40 m/s, the force at the rear undergoes a sharp growth. It is worth noticing that this sharp variation of the loads introduces a significant displacement of the centre of loads during the water entry, which may have an important effect on the aircraft dynamics in a real ditching scenario. The total loads acting in the $z$- and $x$-directions are provided in the lower panels of Figure 17. The data indicate that the $x$-component is much smaller than the $z$-component and is characterized by a small hump occurring about $\tau$= 0.5. Although already visible in terms of the rear and forward components, the total $z$-force is characterized by a large increase with the horizontal speed, starting from the time at which the ventilation starts. Before that time, the positive force in the front part is partly balanced by the negative loads at the rear. When ventilation starts, there is a significant increase in the total loads due to the jump in the pressure values acting on the rear portion of the specimen. A positive contribution in this time interval also derives from the increase in the front force, as a consequence of the growth of the wetted area and to the forward propagation of the spray root, with the associated pressure peak.
By applying to the loads the same scaling proposed for the pressures in (Iafrati, 2016; Spinosa & Iafrati, 2021), the forces can be non-dimensionalized by the factor $\rho\,\left(U^{2}+V^{2}\right)\,S$ where $S$ is the area of the projection of the specimen on the $x-y$ plane. The non-dimensional forces plotted in Figure 18 indicate that the scaling holds for the forward component but not for the rear one, as a consequence of the occurrence of cavitation and ventilation.
Figure 18: Time histories of the forces scaled by the quantity $\rho\,(U^{2}+V^{2})\,S$ for the tests of shape S2, at pitch angle 6 and different horizontal speeds.
3.4 Effect of the pitch angle
The effect of the pitch angle on pressures and loads is somewhat more complicated by the fact that the first contact point H changes with it, and therefore the local geometry, in particular the longitudinal curvature around the first contact point, is different. This feature is highlighted in Figure 19, where the mid-plane sections of the specimen S2 at the three pitch angles of 4, 6 and 8, aligned with respect to the first contact point, are plotted.
Figure 19: Comparison among the midline profiles of the shape S2 at different pitch angles and aligned with respect to the first contact point H.
It is worth noticing that the differences among the three profiles are not so relevant behind H but are very evident in the front part. As the first contact point H moves backwards as the pitch angle increases, the relative positions of the probes with respect to the point of initial contact are not the same, as shown in Figure 20.
Figure 20: Position of the curvature change point C and of the points of first contact H4, H6 and H8, which correspond to the pitch angles of 4, 6 and 8, respectively.
In figure 21 the underwater frames at $\tau$= 0.6 and $\tau$= 0.9 for the different pitch angles are shown, together with the edges of geometric intersection between the specimen inclined at different pitch angles and the still water level at $\tau$=0.6 and $\tau$=0.9.
[…]
Figure 21: Under-water frames comparison at $\tau$= 0.6 and 0.9 for the specimen S2 at $U$=40 m/s and at the pitch angles 4, 6 and 8.
By comparing the wetted area with the geometric intersection areas, a large consistency in found, aside from the presence of the cavitation and ventilation zones at the rear as well as the pile-up and the spray region in front.
The expansion of the cavitation zone to the trailing edge is observed to be faster at higher pitch angles and consequently the onset of ventilation occurs earlier. Such a situation is expected. In fact, as clearly shown in Figure 19 the slope gradient about the point H, while traversing the lower specimen profile from the leading edge to the trailing edge, is more pronounced in the case of higher pitch angles. This results in the development of lower pressures at the rear during the impact phase, thereby facilitating a faster expansion of the cavitation region. In addition, it is observed that at 6 and at 8 and at $\tau$=0.6 the cavitation region has already reached the trailing edge and a ventilation front has already started propagating forwards. Instead, at 4 the shape of the cavitation region has not changed relevantly between $\tau$=0.6 and $\tau$=0.9. Such a situation is similar to the case of the intermediate cavitation regimes discussed in Section 3.3.
In Figure 22(a-f) and in Figure 22(g-l) the time histories of the pressures recorded by the probes located along the midline l1 and line l3 are shown for the groups of rows from $r5$ to $r8$ and from $r1$ to $r4$, respectively.
(a) $p$ midline - Front probes - Pitch 4
(b) $p$ midline - Front probes - Pitch 6
(c) $p$ midline - Front probes - Pitch 8
(d) $p$ line $l3$ Front probes - Pitch 4
(e) $p$ line $l3$ Front probes - Pitch 6
(f) $p$ line $l3$ Front probes - Pitch 8
(g) $p$ midline - Rear probes - Pitch 4
(h) $p$ midline - Rear probes - Pitch 6
(i) $p$ midline - Rear probes - Pitch 8
(j) $p$ line $l3$ Rear probes - Pitch 4
(k) $p$ line $l3$ Rear probes - Pitch 6
(l) $p$ line $l3$ Rear probes - Pitch 8
Figure 22: Front and rear pressure probe time histories for the specimen S2 at $U$=40 m/s and at the pitch angles 4, 6 and 8.
Looking at the pressure time histories in the front part of the specimen in Figure 22(a)-(f), it is observed that the pressure peaks are of the same order of magnitude for all three shapes and are delayed on the line $l3$, as a consequence of the curvature of the spray root region induced by both the body curvature and three-dimensional effects. The effect of the pitch angle is more evident by looking at the pressures values after the peak, which approach an almost constant value, higher for larger pitch angles, in agreement with what found in Iafrati (2016) for the impact of a flat plate. Something similar happens for the pressures recorded by the probes located along the sensor line $l3$.
As for the time histories recorded by the rear pressure probes, shown in Figure 22(g)-(l), it is shown that the drop to the vapour pressure values of all the probes located along the mid-line occurs over a shorter time period going from $\alpha$= 4 to $\alpha$= 6 and then $\alpha$= 8. Such a behaviour also is noticed at sensor line $l3$, where all pressure signals are delayed with respect to the midline due to the transverse curvature of the body. The pressure time histories also confirm that the duration of the cavitation phase is longer at 4, where ventilation starts just at the end of the impact phase, while at both 6 and 8 it starts at about $\tau=0.6$. It also appears that at 6 and 8, when ventilation occurs, the pressure rise to the atmospheric pressure is quite sharp. In contrast, at 4, the pressure growth is much milder. Such a pressure behaviour, in agreement with what discussed above regarding the video frames, is again similar to the intermediate regimes of cavitation, also observed for the shape S3 in Iafrati & Grizzi (2019).
As the pitch angle increases, it is observed that the peak pressure recorded by the P17 probe also increases. At 4 no evident peak appears. An explanation for such a circumstance could be that, as shown in Figure 20, since the probe P17 is located very close to the first contact, some air cushioning effects may enter into play. Air cushioning may induce a local lowering of the free surface, especially for water entry at high speed and at low pitch angles, as discussed for flat plates in Iafrati et al. (2015). Air entrapment may also occur at higher pitch angles during the initial moments of the water impact, which may explain why a positive pressure peak is not recorded by the probe P13, which at 8 pitch is located at the first contact point.
The time histories of the loads at the three pitch angles are shown in Figure 23.
Figure 23: Time histories of the forces at the three pitch angles for the shape S2 at U=40 m/s.
It can be observed that, as expected, the forces at the front increase as the pitch angle increases by about a factor 2 or more, as the pitch angle passes from 4 to 8. The forces at the rear remain very low up to about half of the impact phase. Then, at 4 the negative loads are more evident, whereas at 8 a sudden jump to higher values occurs as a consequence of ventilation.
3.5 Effects of the body curvatures
Another aspect worth of investigation is the role played by the body curvatures. As already detailed in Section 2.2, three different shapes are considered. The specimens S2 and S3 have the same longitudinal profile in the mid-plane but different curvatures in the transverse direction, S3 being characterized by a rather “flat” elliptical cross-section, while S2 by a circular cross-section. The specimen S1B has a somewhat smoother longitudinal profile compared to the others and a circular cross section, but with a smaller radius than S2.
In order to gain insight on the shape on the spray in front and of the cavitation and ventilation phenomena at the rear, the underwater video-frames taken for the different shapes from $\tau$=0 to $\tau$=0.8 are shown in Figure 24. The frames are cropped and aligned with respect to the specimen leading edge, to facilitate the comparison. The edges of the geometric intersection areas with the still water level from $\tau$=0.2 to 0.6 are superimposed to the corresponding pictures.
(a) Shape S1B
(b) Shape S2
(c) Shape S3
Figure 24: Under-water video frames of the specimens S1B, S2 and S3 during the water entry $U$=40 m/s and pitch angle 6 from $\tau$=0 to $\tau=0.8$. For the frames from $\tau$=0.2 to $\tau=0.6$ the edges of the geometric intersection area with the still water level are also plotted in blue. The direction of water entry is from bottom to the top.
By looking at the underwater images, it can be seen that for S1B, due to the smoother longitudinal curvature, the first point of contact is in the centre of the specimen. (see also Figure 4) and, differently from S2 and S3, the wetted area is rather symmetric with respect to the transverse line passing through the first contact point. In the time interval from $\tau=0.2$ to $\tau=0.6$ it is also evident that for S1B the geometric intersection area exhibits analogous symmetry and is evenly smaller than the wetted area in all directions, which implies a rather uniform pile-up all around. Transversally, the width of the wetted area reflects the differences in the cross sections, therefore S1B has the narrowest section, due to the smaller radius of curvature, whereas S3 has the broadest one due, to the elliptical profile.
The difference in the transverse curvatures is also responsible for the differences in the curvature of the forward propagating spray.
It is also evident that in S1B no cavitation region appears during the impact phase, unlike in S2 and S3. As explained in the previous sections, in a $2D+t$ approximation, the longitudinal curvature at the rear of the body transform the water entry problem into a water exit one. As such, the smoother is the longitudinal curvature, the lower is the water exit speed and the lower is the reduction in the pressure.
As observed for the shape S2 in Section 3.2, also for the shape S3 the cavitation region expands in the rear part, whereas the front edge appears to be fixed at approximately the point of curvature change C. This is evident in the frames shown in Figure 24. As already noted in Section 3.3 the differing transverse profile between S2 and S3 does not alter the fact that the same regimes of evolution of the cavitation region occur, and the threshold values of $U$ that delineate them are similar. However, in agreement with the findings of Spinosa et al. (2022) the pressures at the rear in S3 are expected to be averagely lower (more negative), due to the lower three-dimensionality of the flow. This results in the cavitation region having a faster expansion, which also leads to an earlier formation of the ventilation zone for this shape. A more detailed analysis of the backwards expansion of the cavitation region is provided in Section 3.6, where this assertion is verified in a quantitative manner.
Although in a much less evident form, cavitation occurs also for specimen S1B at the same impact conditions but, as shown in Figure 25, at a much later stage starting from $\tau\simeq 1.9$.
Figure 25: Cavitation area observed in the shape S1B after the end of the impact phase for the test at $U$ =40 m/s and pitch angle 6 . The points H and C are also superimposed. The direction of water entry is from bottom to the top.
In this case, the front edge of the cavitation region appears to be situated in close proximity of the first contact point, whereas the rear edge does not exhibit a significant backward expansion in time. This situation is somewhat consistent with the incipient cavitation regime described in Section 3.3 and in Iafrati & Grizzi (2019). However, as it can be seen in Figure 25, at those times the leading edge of the specimen is already below the still water level, and a strong jet is formed at the front, which propagates towards the trailing edge. In this phase the trolley undergoes a sharp deceleration, and therefore the flow conditions are rather uncertain. For this reason the dynamics of the cavitation region for the specimen S1B is not discussed further.
The above considerations are supported by the time histories of pressures along the mid-line and along line $l3$, which are shown in Figure 26. It is worth noting that in the shape S1B, due to the different location of the first contact point, the probes located in front portion do not include those on the row r5 (see again Figure 20).
(a) $p$ midline - Front probes - Shape S1B
(b) $p$ midline - Front probes - Shape S2
(c) $p$ midline - Front probes - Shape S3
(d) $p$ Line l3 - Front probes - Shape S1B
(e) $p$ Line l3 - Front probes - Shape S2
(f) $p$ Line l3 - Front probes - Shape S3
(g) $p$ midline - Rear probes - Shape S1B
(h) $p$ midline - Rear probes - Shape S2
(i) $p$ midline - Rear probes - Shape S3
(j) $p$ Line l3 - Rear probes - Shape S1B
(k) $p$ Line l3 - Rear probes - Shape S2
(l) $p$ Line l3 - Rear probes - Shape S3
Figure 26: Front and rear pressure time histories for the three shapes, at $U$=40 m/s and at a pitch angle of 6.
Looking at the mid-line probes located in the front part of the specimen, apart from some differences in the peak values, the pressures after the peak display a rather similar behaviour. The effect of the different curvature in the transverse direction appears evident when comparing the time delays between probes located across the same sensor row. The largest delays appears in S1B, whereas the the smallest one in S3. This circumstance can be attributed to the larger curvature of the spray root in the specimens with a higher transverse curvature, which allows for a greater possibility of the fluid to escape from the sides. For the same reason, the pressure values recorded by the probes along the line l3 are lower for the specimens with higher transverse curvature, i.e. S1 and S2.
The time histories of the pressure measured by the rear probes for S1B are generally positive. Differently from what happens for S3, where the probes display a peak of small amplitude followed by a rapid descent and a sharp drop to the negative vapour pressure value, for S1B the pressures display a sharp rise to a maximum value and then a very gentle decay, with some probes assuming negative values at about the end of the impact phase.
Comparing the S2 and the S3 pressure time histories at the rear, it can be seen that, presumably due to the lower transverse curvature and the lower local deadrise angle, evident pressure peaks appear only in S3. In the next stage, for both S2 and S3 the pressures at the rear exhibit sharp drops to the vapour pressure value, which is consistent with what shown by the underwater images. The delay between the pressure drops recorded by the probes placed on the same row between lines $l1$ and $l3$ (for example between P17 and P15 or between P13 and P11) is greater in the case of S2, compared to S3. This is a consequence of the larger curvature of the rear edge of the cavitation region, which appears as a D-shape more elongated in the longitudinal direction (see Figure 24(b) and (c)). The ventilation phase starts earlier in S3 than in S2, which confirms that the expansion of the cavitation region is faster for the former shape, as the cavity reaches the trailing edge earlier in time.
The difference between the the pressure fields of S2 and S3 on the bottom can also be visualized when comparing the iso-lines at $\tau$=0.3, 0.6 and 0.8 plotted for the shape S3 in Figure 27 with those of S2 shown above in Figure 9.
(a) $\tau$= 0.3
(b) $\tau$= 0.6
(c) $\tau$= 0.8
Figure 27: Pressure iso-lines in the rear part of the specimen overlapped to the underwater video frames for the test of the shape S3 at U=40 m/s and at $\alpha=6^{\circ}$ at $\tau$=0.3, 0.6 and 0.8. The symmetry about the mid-line is exploited.
Firstly, it can be observed that, due to the “flatter transverse profile” the pressure at the front edge of the cavitation region rises more sharply and uniformly in the transverse direction for the shape S3, compared to S2. At $\tau$=0.3 the shape of the cavitation area in S3 is still in agreement with the pressure iso-lines, but it has a less elongated shape compared to S2. However, as the cavitation region expands faster in S3, at $\tau$=0.6 the rear edge has already reached the trailing edge and the ventilation front has already moved forwards, up to the point C. At $\tau$=0.8, the development of the ventilation zone has further evolved.
Ventilation in S3 occurs in between $\tau$=0.5 and $\tau$=0.6. In Figure 28 the evolution of the process is visualised at six time instants within this non-dimensional time interval.
(a) Shape S2 - $\tau$=0.50
(b) Shape S2 - $\tau$=0.52
(c) Shape S2 - $\tau$=0.54
(d) Shape S2 - $\tau$=0.56
(e) Shape S2 - $\tau$=0.58
(f) Shape S2 - $\tau$=0.6
[…]
The mechanism of formation and development of the ventilation front in S3 is similar to S2, with a curved front starting from the centre at the trailing edge, which evolves very shortly into a straight one and reaches the point C.
The comparison in terms of the $z$-force time histories for the three shapes is shown in Figure 29.
Figure 29: Time histories of the forces for the three shapes at $U$=40 m/s and pitch angle 6.
The data indicates that the load measured at the front is characterised by an increasing trend for all three shapes. The growth rate is higher, as the transverse curvature decreases, which is consistent with the pressure behaviour discussed above. The load at the rear exhibit a slightly decreasing trend for S1B, whereas for both S2 and S3 an initial decreasing trend is observed, associated with the onset of cavitation, followed by a rather sharp rise, related to the ventilation phase. As ventilation starts earlier for shape S3, the sharp rise occurs earlier, consistently with the pressure time histories. No substantial differences are found for the $x$-component of the force for the three specimens, which is always very small compared to the $z$-component.
3.6 Cavitation zone propagation velocity
The underwater videos and pressure data can be exploited to derive more precise information on the dynamics of the cavitation region, not only from the qualitative point of view, but also from the quantitative point of view. The occurrence of cavitation exerts significant effects in terms of both load amplitude and moments, and thus the centre of loads. It is therefore of great importance to analyse its evolution in more detail, and to evaluate, on a quantitative basis, the influence of the body curvatures upon it, especially the transverse curvature.
For this purpose a propagation velocity of the cavitation region, denoted as $U_{\textrm{cav}}$, is introduced, which is the velocity at which the rear edge of the cavitation zone moves backwards in the trolley reference frame. This propagation velocity can be retrieved either from the analysis of the underwater video frames or from the pressure signals, as follows. Starting from the underwater images the edge of the cavitation region is identified and the furthest point at the back is marked for all frames, which are separated by a constant time interval. The position of the rear edge in the pictures can be determined within $\pm$ 2 pixels, which makes the calculation of the propagation velocity rather accurate. The propagation velocity is computed as an average velocity from the instant in which the cavitation region clearly appears in the frames up to the instant at which it reaches its maximum extension, i.e. when the region reaches a steady size in the incipient cavitation regime or as soon as the region reaches the trailing edge in the cavitation-ventilation regime. The propagation velocity of the cavitation region can also be determined from the time histories of the pressure measured by the different probes. In fact, the passage of the cavitation front corresponds to the sharp drop to the vapour pressure value. Hence, being the position of the probes on the specimen known, the velocity can be determined from the time delay between the pressure drops recorded by different probes. Since the raw pressure signals, as observed above, are often very noisy and display significant spikes, to facilitate the estimation of the time delays, a de-noising procedure has been developed in MATLAB©, which includes in sequence a Continuous Wavelet Transform - based filter, a conventional low-pass filter and a moving average. Hence, the time delay is calculated through a cross-correlation of the denoised signals, with the objective of identifying the time shift that optimises the degree of overlap between the drop portions. The process is illustrated in detail in Figure 30.
(a) Signal De-Noising
(b) Delay estimation
Figure 30: Comparison of (a) two sample signals recorded by the pressure probes P4 and P9 before and after the de-noising procedure and (b) overlapping of the signal P9 and the signal P4, which is the one recorded by P4 and translated in time for the delay estimation.
The estimated propagation velocities of the cavitation region $U_{\textrm{cav}}$, based on the image analysis and on the pressure time histories, are plotted in Figure 31 for the shapes S2 and S3 at 6, velocity ratio $V/U$=0.0375, and different horizontal speeds. It is worth noticing that in the graph supplementary test conditions with different horizontal velocities are considered, in addition to those listed in Table 2. $U_{\textrm{cav}}$ is plotted as a function of the horizontal velocity $U$, and linear fits of the data derived from the image analysis are also plotted, which both have a coefficient of determination $R^{2}$ very close to 1. It is worth noticing that the propagation velocity computed by using the pressure signals are very much in line with those derived from the image analysis.
Figure 31: Propagation velocities of the cavitation area for the three shapes and as a function of $U$, using the method bases on image analysis (circles) and on pressure data (triangles). Linear trend lines are added using the results of the image analysis method.
The above results indicate that $U_{\textrm{cav}}$ is directly proportional to the horizontal speed U and that the coefficient of proportionality is between 1.77 and 1.91 for the shapes S2 and S3, respectively. There is a vertical offset between the two trends, with $U_{\textrm{cav}}$ being higher for S3 compared to S2, as a consequence of the lower transverse curvature that reduces the three-dimensional effects on the resulting flow, consistently with the considerations presented in Section 3.5.
3.7 Wetted area evolution
In the pioneering works of Von Karman (1929) and Wagner (1932) on vertical water entry, the growth of the wetted area is related to the time variation of the added mass and, therefore, to the impact loads. The primary distinction between the two theories is that, whereas von Karman’s theory considers solely the geometric intersection of the body with the free, the Wagner method also accounts for the water pile-up at the spray. In the work of Iafrati (2016) on the water entry of flat plates with a high horizontal speed, it was observed that the value of the pressure peak and the propagation of the spray root during the impact phase are related. Consequently, the evolution of the wetted area and the loads are also subject to a similar relationship. In light of the aforementioned discussion, it can be reasonably assumed that a comparable analogy can be expected in the front part of double-curvature specimens, where the pressures are found to be positive in all areas. Conversely, at the rear, the occurrence of negative pressures and, consequently, of cavitation and ventilation renders such an analogy invalid. In light of the aforementioned considerations, it is deemed insightful to conduct a more detailed investigation into the evolution of the wetted area in relation to the geometric intersection area.
Looking again at Figure 24, referring to the water entry of the three shapes at U=40 m/s, $\alpha$=6 and V/U=0.0375, as mentioned already in Section 3.2, the wetted areas are identifiable as the darker regions at the centre of the specimen. Inside the wetted area the brighter zone, when present, is the cavitation region. It is observed that, due to the pile-up effect, for S1B and S2 the actual wetted area is broader than the geometric intersection area with the undisturbed free surface at $\tau$ in between 0.2 and $\tau$= 0.6. This is also the case for the shape S3 up to $\tau$= 0.3. However, for S3 from a value of $\tau$ between 0.4 and 0.5, if the pile-up effect persists at the front, at the rear of the specimen, in addition to the cavitation/ventilation zone at the middle, some bright zones emerge at the sides. These zones are most likely caused by ventilation originating from the sides of the specimen and developing as a result of the low pressures in the middle. Two dark narrow regions of wetted area are still present between the cavitation/ventilation zones in the middle and the ventilation zones at the sides. In particular, it is observed that for S3 the wetted area becomes smaller than the geometrical intersection area at a non-dimensional time $\tau$ between 0.4 and 0.5. Such an effect becomes more evident at $\tau$= 0.6. A similar effect was also observed in Spinosa et al. (2022), in which a shrinking of the wetted area with respect to the geometric intersection area is observed in the planing motion of the whole fuselages S2 and S3, with the effect more enhanced for S3, as a consequence of the lower pressures at the rear.
From the quantitative point of view it is possible to estimate the wetted area from the underwater video frames and to compare it with the geometrical intersection area. This is performed for the three shapes, using an image processing routine developed in MATLAB©. As shown in Figure 32, which refers to shape S3 at $\tau$=0.2 and 0.6, three Region Of Interests (ROIs) are manually selected and their areas are computed, first as pixel counts and then as the physical area in m2, by multiplying the pixel count for the pixel area derived from an image calibration. In particular, the defined ROIs are the front and rear wetted surface, which are delimited by the $x$-coordinate of the pressure probe row $r4$, the front and rear part of the geometric intersection area, still separated by the row $r4$, and finally the cavitation area. It is worth noticing that using this routine, the wetted areas are not exactly estimated rather their projection on the plane $X-Y$, parallel to the ground.
(a) Wetted Area Front/Rear $\tau$=0.3
(b) Geometric Intersection Front/Rear $\tau$=0.3
(c) Cavitation Area $\tau$=0.3
(d) Wetted Area Front/Rear $\tau$=0.6
(e) Geometric Intersection Front/Rear $\tau$=0.6
(f) Cavitation Area $\tau$=0.6
Figure 32: Definition of the ROI for (a,d) the wetted areas in front and at the rear of the specimen; (b,e) geometrical intersection area at the front and at the rear; (c,f) cavitation area for the shape S3 at U=40 m/s, pitch angle 6 at $\tau$=0.3 (top) and 0.6 (bottom).
A comparison between the values of front/rear wetted areas and of the geometrical intersection from $\tau$=0.2 to 0.6 (see also Figure 24) is shown in Figure 33 for the three shapes The values of the cavitation areas are plotted in Figure 33(g) and (h) for the shapes S2 and S3 respectively.
(a) Front part S1B
(b) Front part S2
(c) Front part S3
(d) Rear part S1B
(e) Rear part S2
(f) Rear part S3
(g) Cavitation area S2
(h) Cavitation area S3
Figure 33: Variation in time wetted areas compared with the geometrical intersection area for the specimen S1B(a,d), S2 (b,e) and S3(c,f) at pitch angle 6 at U=40 m/s from $\tau$= 0.2 to 0.6. In (g,h) the cavitation area as a function of $\tau$ is shown for the shapes S2 and S3.
[…]
At the rear, in the case of specimen S1B both the geometric intersection and the wetted areas grow in time, the latter at a slightly higher rate. In the case of S2 and S3, the values of the geometric intersection area display a monotonic trend, although in the case of S3 the growth rate exhibits a reduction after $\tau=0.4$. The wetted area grows up to $\tau=0.5$ and $\tau=0.4$ for S2 and S3, respectively. For S2 the wetted surface remains almost constant afterwards, as it reaches the edges of the specimen. In the case of S3 the occurrence of ventilation at the sides causes a shrinking of the wetted area, which progressively approaches the value of the cavitation area. For both S2 and S3 the cavitation area grows almost linearly in time, although for S3 it approaches an almost constant value after $\tau$=0.5, once the cavitation zone has reached the trailing edge of the specimen, and the rear ventilation starts.
4 Conclusions and Future Work
In this paper the water entry at high horizontal speed of rigid specimens with double curvature has been investigated, based on experimental data. The effects of the horizontal speed and of the pitch angle are analysed, keeping constant the vertical-to-horizontal velocity ratio. Three different shapes with different transverse and longitudinal curvatures have been considered.
It has been observed that the presence of a longitudinal curvature plays a key role, marking an important distinction with respect to the water entry of flat plates at similar test conditions (Iafrati, 2016). Indeed, the longitudinal curvature, depending on velocity ratio and on the pitch angles, can lead to the occurrence of negative pressures (relative to atmospheric pressure) at the rear, which can reach the vapour pressure value. This, in turn, can result in the formation of a cavitation region. Furthermore, as the region of low pressure at the rear may expands significantly during the water entry, it can favour air entrainment either from the trailing edge or from the sides, leading to ventilation. It has been shown that the transverse curvature does not influence much the cavitation modalities, however it has an effect on the loads and on the dynamics of the cavitation region, due to the increased possibility for the fluid to escape from the side, in case of higher curvature.
The pitch angle is found to affect the pressures and the loads in the front part of the body, the loads being higher for higher pitch angles, consistently with what found for the impact of a flat plate in Iafrati (2016). In addition, given the change in the gradient slope in the rear part, the modalities of cavitation may change. At the higher pitch angle a cavitation-ventilation regime is observed, whereas at the lower pitch angles the ventilation region remains confined close to the trailing edge during the impact phase.
Based on the underwater video frames and on the pressure measurements, more detailed information on the dynamics of the cavitation region and of the wetted area can be retrieved. It is observed that the propagation velocity of the cavitation region increases linearly with the horizontal speed both for S2 and S3. Although the slope is approximately equivalent in both case, an offset exists between the two trends, being the propagation speed higher for S3, as a consequence of the lower transverse curvature.
Starting from the underwater images, quantitative estimates of the wetted area, of the the geometric intersection area with the still water level and of the cavitation area can be retrieved. A distinction is made between what happens ahead and behind the point of curvature change. The data indicate that for the shape S1B, the growth rate of all areas is lower, and so is the difference between the wetted area and the geometric intersection area, compared to the specimens characterised by higher longitudinal curvatures. For shape S3 a ventilation phenomenon from the sides has been also observed, occurring at the rear, as a consequence of the low pressures developing beneath. The occurrence of such ventilation zones causes a shrinking of the wetted area at the rear, which becomes smaller than the corresponding geometric intersection area.
In addition to contributing to a deeper understanding of the physical phenomena occurring during the water entry at high horizontal speed and of the effects induced by changes in the body shape, the present paper provides a rather wide and detailed dataset, which could be used for an accurate validation of numerical approaches for the simulation of aircraft ditching or other water entry problems. Future work will focus on developing routine-based pattern recognition and machine learning algorithms to automate the selection and tracking in time of the wetted area, cavitation areas and the following ventilation front in underwater videos, which can be improved by also exploiting the information coming from the pressure probes.
Acknowledgements
The authors also wish to thank: Flavio Olivieri, the technical supervisor of the High Speed Ditching Facility, who is also responsible for its maintenance. He also participated in the the execution of all the experiments; Fabio Carta, the project manager who dealt with the design of the specimens and with the installation of the instrumentation on board of them and on the facility; Mauro Sale who provided technical support during the execution of the experiments; Alessandro Del Buono was of value in providing assistance with the interpretation of the results.
Funding
This project has been partly funded from the European Union’s Horizon 2020 Research and Innovation Programme under Grant Agreement No. 724139 (H2020-SARAH: increased SAfety & Robust certification for ditching of Aircrafts & Helicopters).
Declaration of interests
The authors report no conflict of interest.
Author contributions
Alessandro Iafrati and Silvano Grizzi designed the facility, installed the experimental setup and performed the experiments. Emanuele Spinosa and Silvano Grizzi carried out the data analysis. Emanuele Spinosa organized and prepared the article drafts. Alessandro Iafrati revised the drafts and supervised the article preparation.